OpenStudy (anonymous):

A football is kicked from the ground at 39.5m/s at an angle of 40.0degrees above the horizontal. At what two times does it reach an altitude of 12.0m

OpenStudy (jdoe0001):

hmmm is there an equation you're meant to use?

OpenStudy (anonymous):

this is physics. he's supposed to use known equations. what exactly is you issue with the problem? @apink<3 ?

OpenStudy (anonymous):

its projectiles and the equations are basically the big 4 and v = d/t

OpenStudy (anonymous):

im suppose to find the two times of the ball at a altitude of 12m so basically the time up and down

OpenStudy (anonymous):

Use: \[\Delta s = V_0 t + \frac{ 1 }{ 2 }at^2\] since the forces/accelerations in the x and y directions are independent, and we only care about the y direction, we use this equation for the y direction: \[\Delta y = V_0 (\sin \theta )t + \frac{ 1 }{ 2 } g t^2\] solving for t as you would a polynomial (quadratic eqn)